OpenStudy (anonymous):

Rae earns $60 for playing music on Saturday nights at the local coffee house. She makes an average of $5.25 in tips per hour. Write the function of Rae's earnings and solve for how much she would make after 4 hours. y = 5.25x + 60; $65.25 y = 60x + 5.25; $85.75 y = 5.25x + 60; $81.00 y = 60x + 5.25; $185.25

OpenStudy (anonymous):

y=5.25(4)+60 y=21+60 y=81